Tony is 20 years older than his sister Danielle. In 16 years he will be twice as old as her. how old is each of them now?
Thanks

Answer by jorel1380(3719)

Tony=Danielle+20
T+16=2(D+16)
D+20+16=2D+32
D+36=2D+32
D=4
T=24
Tony is 24. Danielle is 4..